OpenStudy (anonymous):

What is the anti derivative of (3/ (1+9s^2))?? need help

OpenStudy (mathteacher1729):

\[3\int \frac{dx}{1+(3x)^2}\] This should look familiar... What inegral has the form \[\frac{du}{1+u^2}\] ?

OpenStudy (anonymous):

arctan